# Value-Based Large Language Model Agent Simulation for Mutual Evaluation of Trust and Interpersonal Closeness

### v1

@@ Results
This preliminary experiment evaluated the value controllability of LLMs via prompting. We quantified the LLM responses to instructions to adopt specific values using a perspective controllability score [8]. We assessed this controllability across several factors: prompt design (system vs. user prompt; second- vs. third-person prompt; inclusion vs. exclusion of value definitions), value dimensions (the ten basic and four higher-order values from Schwartz’s theory of basic values), and language (English and Japanese). The experimental procedure is described in detail in the Methods section.
Tables 1 and 2 summarize the measured perspective controllability scores for the English and Japanese languages, respectively. In English (Table 1), the results indicate that Gemini demonstrated the highest controllability. Specifically, it achieved the highest score for basic values under the “system prompt, second-person, with definition” condition, and for higher-order values under the “user prompt, second-person, with definition” condition. Similarly, in Japanese (Table 2), Gemini exhibited the highest controllability. The highest score for basic values was obtained under the “system prompt, second-person, with definition” condition, whereas for higher-order values, the “system prompt, second-person, without definition” condition yielded the top score.
@@
Prompt Person Definition Basic Higher-order Basic Higher-order Basic Higher-order
System 2 Yes 0.8216 0.8766 0.5931 0.6836
No 0.7315 0.8768 0.5772 0.6823
3 Yes 0.7678 0.8432 0.6251 0.6119 0.6937
No 0.7123 0.8361 0.6752 0.7870 0.6014 0.7038
User 2 Yes 0.8142 0.8838 0.6552 0.7483
No 0.7692 0.8553 0.6485 0.7463
3 Yes 0.6731 0.7618
No 0.6816 0.7719
Average 0.7694 0.8620 0.6502 0.7870 0.6303 0.7240
Table 1: Perspective controllability scores in English. The highest score in each column is shown in bold.
@@
Prompt Person Definition Basic Higher-order Basic Higher-order Basic Higher-order
System 2 Yes 0.6960 0.7796 0.4729 0.6071
No 0.6237 0.7824
3 Yes 0.4784 0.5949 0.4982 0.6643
No 0.3442 0.5733 0.5166 0.5929 0.5097 0.6811
User 2 Yes 0.6959 0.7647 0.6782
No 0.6269 0.7636
@@ 3 Yes
Average 0.5775 0.7098 0.5166 0.5929 0.4936 0.6577
Table 2: Perspective controllability scores in Japanese. The highest score in each column is shown in bold.
@@
The mutual evaluation scores under all conditions in this experiment are shown in Figures 2 and 3 for English and Figures 4 and 5 for Japanese. The means and standard deviations of the data under each condition are listed in Table 3. These results indicate a tendency toward higher trust and interpersonal closeness scores between agents with greater value similarity. For instance, in Figure 2(a), the agent pairings that received the highest trust score of 5.0 were those with highly similar values, such as universalism–universalism and tradition evaluating conformity. Conversely, one of the pairings with the lowest score of 1.0 (Figure 2(a)) was hedonism evaluating conformity; these two values are in opposing positions within Schwartz’s theory of basic values.
@@
We categorized the similarity of values between agents into three levels (high, medium, and low) based on Schwartz’s theory of basic values and calculated the average mutual evaluation scores for each level to quantitatively evaluate these visual observations (Tables 4 and 5). As summarized in Tables 4 and 5, a trend was observed wherein a higher value similarity corresponded to higher scores for trust and interpersonal closeness, regardless of the language or dialogue task.
[…]
(a) Trust, basic values.
(b) IOS, basic values.
(c) Trust, higher-order values.
(d) IOS, higher-order values.
Figure 2: Results of mutual evaluation (collaborative decision-making about housing, English).
(a) Trust, basic values.
(b) IOS, basic values.
(c) Trust, higher-order values.
(d) IOS, higher-order values.
Figure 3: Results of mutual evaluation (dialogue about hobbies, English).
(a) Trust, basic values.
(b) IOS, basic values.
(c) Trust, higher-order values.
(d) IOS, higher-order values.
Figure 4: Results of mutual evaluation (collaborative decision-making about housing, Japanese).
(a) Trust, basic values.
(b) IOS, basic values.
(c) Trust, higher-order values.
(d) IOS, higher-order values.
Figure 5: Results of mutual evaluation (dialogue about hobbies, Japanese).
English Japanese
Housing Hobbies Housing Hobbies
Metric Basic Higher-order Basic Higher-order Basic Higher-order Basic Higher-order
trust
$3.151\pm$
1.516
$3.655\pm$
1.277
$2.522\pm$
1.497
$3.174\pm$
1.471
$3.721\pm$
0.971
$3.651\pm$
1.108
$3.457\pm$
0.900
$3.164\pm$
1.187
IOS
$5.404\pm$
1.310
$5.725\pm$
0.862
$3.780\pm$
2.323
$4.950\pm$
1.904
$5.829\pm$
0.519
$5.835\pm$
0.538
$5.617\pm$
0.956
$5.465\pm$
1.260
Table 3: Mean and standard deviation in each setting.
@@
Table 4: Relationship between value similarity and mutual evaluation results (Trust).
@@
Table 5: Relationship between value similarity and mutual evaluation results (IOS).
@@
[…]
@@ Discussion
The results of the preliminary experiment on value controllability revealed that the perspective controllability scores varied depending on the target language model, prompt format, and language (Tables 1 and 2). First, a comparison of the performance of the models revealed that Gemini achieved the highest average scores in both English and Japanese. This suggests that, Gemini has a better ability to generate responses aligned with a specified perspective than the other models under the conditions of this study. Next, we observed two trends regarding the influence of the prompt format. First, controlling the models based on the ten basic values was consistently more challenging than controlling the models based on the four higher-order values. It is conceivable that, because the latter represents broader and more abstract concepts, the models can more easily grasp them as general directives, leading to improved controllability. Second, the effects of grammatical persons varied among the models. For Gemini, second-person prompts yielded greater controllability than third-person prompts. By contrast, GPT exhibited the opposite trend, in which third-person prompts tended to result in higher controllability. Furthermore, a significant disparity was observed between languages; scores in English consistently surpassed those in Japanese across all models and conditions. This is likely attributable to the fact that the training data of several large language models are predominantly in English. Consequently, we infer that the models’ comprehension and adherence to instructions were higher for English than for Japanese. The existence of this linguistic bias poses a significant challenge in applying LLMs to value-based systems in non-Anglophone contexts.
@@
A further comparison between the English and Japanese results revealed a notable difference: whereas scores were stably calculated in the English experiments, the Japanese experiments exhibited a significant number of missing values, which corresponded to response refusals by the models (Tables 1 and 2). Communication in Japan is often discussed as a classic example of a high-context culture [17], and recent studies have indicated that LLMs reflect the social biases present in training data [18]. Based on these findings, it is conceivable that LLMs learn not only vocabulary and grammatical structures but also statistically internalize communication styles rooted in a high-context culture, such as avoiding direct assertions and circumventing conflict, during their training on Japanese text data. Therefore, we infer that achieving value control in the Japanese context requires sophisticated prompt engineering that goes beyond merely presenting values. Such prompts must also signal to the model the meta-level situational judgment such that “an explicit and assertive evaluation is necessary in this context.” Thus, our findings suggest that the value control of LLMs is a localization challenge that necessitates a deep understanding of the linguistic and cultural context in which the model operates.
@@
Subsequently, the results of the main experiment are discussed. As summarized in Tables 4 and 5, a higher similarity in values, based on Schwartz’s theory of basic values, was associated with higher mutual trust ratings and a greater sense of interpersonal closeness. This finding is consistent with a broad range of previous studies demonstrating the importance of value similarity in trust and relationship-building [19, 12, 11]. Therefore, our results can be interpreted as successful replication of these findings within the computational framework of mutual evaluation. Furthermore, as summarized in Table 6, a positive correlation was observed between the evaluation patterns of the ten basic values and the four higher-order values under most conditions. This finding suggests that Schwartz’s theory [6, 7] was also reproduced within the mutual evaluation framework employed in this study.
@@
Differences in the mutual evaluations based on the topic of interaction were also observed. As summarized in Table 3, evaluations of trust and interpersonal closeness tended to be higher in the cooperative decision-making task regarding housing than in the dialogue regarding hobbies. This finding can be explained by the relationship development process. Interpersonal relationships deepen from the initial stage of simple self-disclosure to a more advanced stage of sharing fundamental values, and the depth of the relationship is linked to trust [21, 22]. In the present study, dialogue about hobbies was assumed to mimic the superficial level of interaction characteristic of the initial stages of a relationship, whereas the cooperative decision-making task was considered to represent interaction at a deeper level, involving the sharing of values. Therefore, the difference in evaluation scores observed in this experiment suggests that topic of dialogue can simulate the process of relationship development.
@@
Furthermore, differences in the mutual evaluations based on language were observed. As summarized in Table 3, the tendency to have higher mean scores and smaller standard deviations for mutual evaluations was observed in Japanese than in English. This trend suggests that the simulation reflects the tendency of the Japanese culture to emphasize harmony and relationships with others [13]. Moreover, the correlation analysis summarized in Table 6 reveals lower correlation coefficients for Japanese than for English. Specifically, no significant correlation was found between the evaluation patterns of basic and higher-order values in cooperative decision-making regarding housing tasks. This may indicate that the structure of values depends on cultural context [6, 13]. Thus, the deviation in the pattern observed under Japanese conditions in this study may reflect this cultural dependency.
@@
Finally, we address the limitations of this study and outline directions for future research. First, a key future task will involve constructing a model that considers the interactions among multiple values. In the present experiment, the value provided to the LLM in the prompt was singular; therefore, the influence of other coexisting values could not be considered. However, humans often hold multiple values and make decisions based on trade-offs between them [23]. Therefore, modeling this value plurality is expected to enable a more realistic value simulation. Second, we propose the development of a more comprehensive agent model that integrates factors beyond values. Human decision-making is influenced not only by values but also by various cognitive and affective factors, such as momentary emotions and the inference of the intentions of others [24]. Incorporating these elements into the agent architecture is likely to enable more human-like simulations under realistic conditions.

### v2

@@ Results
This preliminary experiment evaluated the value controllability of LLMs via prompting. We quantified the LLM responses to instructions to adopt specific values using a perspective controllability score, which indicates the extent to which the model expresses the value or personal trait corresponding to the prompted perspective more strongly than other traits [8]. Controllability was assessed across each value dimension (the ten basic and four higher-order values in Schwartz’s value theory), with parameters including prompt design (system or user prompt, second-person or third-person prompt, inclusion or exclusion of value definitions) and language (English and Japanese). The system prompt refers to the instruction primarily used to set the model’s overall behavior and role, whereas the user prompt is the input provided by the user during the interaction. These are standard functionalities provided in the APIs of large language models (LLMs). In the second-person prompt, the model is addressed directly using “you.” In the third-person prompt, the prompt starts with the phrase “The following are answers from a person …”. The experimental procedure is described in detail in the Methods section.
The measured perspective controllability scores are presented in Table 1 for the ten basic values and in Table 2 for the four higher-order values. For the ten basic values, the highest controllability was observed under the combination of “Gemini, system prompt, second-person, with definition,” in both English and Japanese. For the four higher-order values, the highest controllability was observed under “Gemini, user prompt, second-person, with definition” in English, and under “Gemini, system prompt, second-person, without definition” in Japanese.
Following prior work [8], we conducted Welch’s t-tests with Bonferroni correction to compare controllability across models. Using the prompt design that yielded the highest controllability for each LLM, we compared Gemini—the model with the highest controllability—against Claude and GPT. For the ten basic values, Gemini exhibited significantly higher controllability than both Claude (t(62.76) = 136.19, p < .001) and GPT (t(57.81) = 106.05, p < .001) in English, as well as significantly higher controllability than Claude (t(97.31) = 52.85, p < .001) and GPT (t(78.52) = 40.62, p < .001) in Japanese. For the four higher-order values, Gemini again outperformed Claude (t(68.32) = 49.19, p < .001) and GPT (t(78.84) = 69.37, p < .001) in English, and also showed significantly higher controllability than Claude (t(61.14) = 76.06, p < .001) and GPT (t(63.64) = 44.31, p < .001) in Japanese. Furthermore, to compare controllability between languages, we conducted Welch’s t-tests with Bonferroni correction using the prompt design and LLM that achieved the highest controllability in each language. There was a significant difference between English and Japanese for both the ten basic values, t(51.69) = 53.90, p < .001, and the four higher-order values, t(97.94) = 87.13, p < .001.
@@
Prompt Person Definition English Japanese English Japanese English Japanese
System 2 Yes
$\textbf{0.8216}\pm$
0.0027
$\textbf{0.6961}\pm$
0.0161
$0.5931\pm$
0.0048
$0.4729\pm$
0.0146
No
$0.7315\pm$
0.0089
$0.6237\pm$
0.0037
$0.5772\pm$
0.0078
3 Yes
$0.7678\pm$
0.0029
$0.4784\pm$
0.0168
$0.6251\pm$
0.0063
$0.6119\pm$
0.0039
$0.4982\pm$
0.0195
No
$0.7123\pm$
0.0183
$0.3442\pm$
0.0180
$\textbf{0.6752}\pm$
0.0070
$\textbf{0.5166}\pm$
0.0175
$0.6014\pm$
0.0076
$\textbf{0.5097}\pm$
0.0278
User 2 Yes
$0.8142\pm$
0.0049
$0.6959\pm$
0.0120
$0.6552\pm$
0.0051
No
$0.7692\pm$
0.0074
$0.6269\pm$
0.0159
$0.6485\pm$
0.0077
3 Yes
$0.6731\pm$
0.0059
No
$\textbf{0.6816}\pm$
0.0088
Table 1: Perspective controllability scores in the ten basic values. The table reports the mean and standard deviation of the scores obtained from 50 runs. The highest value in each column is shown in bold.
@@ 3 Yes
Prompt Person Definition English Japanese English Japanese English Japanese
System 2 Yes
$0.8766\pm$
0.0035
$0.7796\pm$
0.0076
$0.6836\pm$
0.0086
$0.6071\pm$
0.0281
No
$0.8768\pm$
0.0090
$\textbf{0.7824}\pm$
0.0058
$0.6823\pm$
0.0089
@@
$0.8361\pm$
0.0062
$0.5733\pm$
0.0122
$\textbf{0.7870}\pm$
0.0125
$\textbf{0.5929}\pm$
0.0164
$0.7038\pm$
0.0090
$\textbf{0.6811}\pm$
0.0149
User 2 Yes
$\textbf{0.8838}\pm$
0.0057
$0.7647\pm$
0.0128
$0.7483\pm$
0.0090
$0.6782\pm$
0.0193
No
$0.8553\pm$
0.0079
$0.7636\pm$
0.0125
$0.7463\pm$
0.0094
3 Yes
$0.7618\pm$
0.0074
No
$0.5903\pm$
0.0189
$\textbf{0.7719}\pm$
0.0098
Table 2: Perspective controllability scores in the 4 higher-order values. The table reports the mean and standard deviation of the scores obtained from 50 runs. The highest value in each column is shown in bold.
@@
The mutual evaluation scores under all conditions in this experiment are shown in Figures 2 and 3 for English and Figures 4 and 5 for Japanese. For each condition, we conducted 10 independent dialogue runs and report the mean and standard deviations of mutual evaluation score. As indicated by the row and column labels in Figures 25, the matrices also indicate which agent evaluates which. In the diagonal cells, the evaluator and the evaluated share the same value; therefore, the cell means are based on twice as many data points as the off-diagonal cells. These results indicate a tendency toward higher trust and interpersonal closeness scores between agents with greater value similarity. For instance, in Figure 2(a), the agent pairings that received the highest trust score of 5.0 were those with highly similar values, such as universalism–universalism and tradition evaluating conformity. Conversely, one of the pairings with the lowest score of 1.0 (Figure 2(a)) was hedonism evaluating conformity; these two values are in opposing positions within Schwartz’s theory of basic values.
@@
We categorized the similarity of values between agents into three levels (high, medium, and low) based on Schwartz’s theory of basic values and calculated the average mutual evaluation scores for each level to quantitatively evaluate these visual observations (Tables 3 and 5). To test whether evaluation scores increased monotonically with the degree of value similarity, Jonckheere–Terpstra tests were conducted separately for each language (English/Japanese), dialogue task (Housing/Hobbies), and value classification (basic values/higher-order values). Across all patterns, the results revealed significant positive trends, indicating that greater value similarity was associated with higher ratings of both trust and IOS (all p < .001; see Tables 4 and 6 for details).
We conducted two-way ANOVAs with language and topic as factors for each evaluation metric (trust / IOS) and each value classification (basic / higher-order) to statistically examine the effects of language and topic on evaluation scores. Across all analyses, a consistent main effect of topic was observed: dialogues on the topic of “housing” yielded higher trust and IOS scores than those on “hobbies” (Trust–basic: $F(1,4396)=139.21$, $p<.001$; Trust–higher-order: $F(1,796)=29.15$, $p<.001$; IOS–basic: $F(1,4396)=446.60$, $p<.001$; IOS–higher-order: $F(1,796)=41.99$, $p<.001$). Regarding the effect of language, significantly higher scores were obtained in Japanese than in English for Trust–basic, IOS–basic, and IOS–higher-order (Trust–basic: $F(1,4396)=396.38$, $p<.001$; IOS–basic: $F(1,4396)=678.73$, $p<.001$; IOS–higher-order: $F(1,796)=12.51$, $p<.001$). In addition, the topic $\times$ language interaction was significant for Trust–basic ($F(1,4396)=23.20$, $p<.001$), IOS–basic ($F(1,4396)=264.23$, $p<.001$), and IOS–higher-order ($F(1,796)=5.25$, $p=.022$). By contrast, for Trust–higher-order, neither the main effect of language ($F(1,796)=0.01$, $p=.938$) nor the interaction ($F(1,796)=0.001$, $p=.974$) reached significance.
Furthermore, we performed a correlation analysis to investigate the relationship between the evaluation patterns at basic and higher-order values (Table 7). The results demonstrated a statistically significant positive correlation in most cases. However, no significant correlation was found for interpersonal closeness in collaborative decision-making about the housing task in Japanese ($r=.2960$, $p=.2657$).
(a) Trust, basic values (M = 3.151, SD = 1.516).
(b) IOS, basic values (M = 5.404, SD = 1.310).
(c) Trust, higher-order values (M = 3.655, SD = 1.277).
(d) IOS, higher-order values (M = 5.725, SD = 0.862).
Figure 2: Results of mutual evaluation (collaborative decision-making about housing, English). Each cell shows the mean score across 10 independent runs, with the standard deviation in parentheses. Rows represent the evaluator agent’s value, and columns represent the evaluated agent’s value.
(a) Trust, basic values (M = 2.522, SD = 1.497).
(b) IOS, basic values (M = 3.780, SD = 2.323).
(c) Trust, higher-order values (M = 3.174, SD = 1.471).
(d) IOS, higher-order values (M = 4.950, SD = 1.904).
Figure 3: Results of mutual evaluation (dialogue about hobbies, English). Each cell shows the mean score across 10 independent runs, with the standard deviation in parentheses. Rows represent the evaluator agent’s value, and columns represent the evaluated agent’s value.
(a) Trust, basic values (M = 3.721, SD = 0.971).
(b) IOS, basic values (M = 5.829, SD = 0.519).
(c) Trust, higher-order values (M = 3.651, SD = 1.108).
(d) IOS, higher-order values (M = 5.835, SD = 0.538).
Figure 4: Results of mutual evaluation (collaborative decision-making about housing, Japanese). Each cell shows the mean score across 10 independent runs, with the standard deviation in parentheses. Rows represent the evaluator agent’s value, and columns represent the evaluated agent’s value.
(a) Trust, basic values (M = 3.457, SD = 0.900).
(b) IOS, basic values (M = 5.617, SD = 0.956).
(c) Trust, higher-order values (M = 3.164, SD = 1.187).
(d) IOS, higher-order values (M = 5.465, SD = 1.260).
Figure 5: Results of mutual evaluation (dialogue about hobbies, Japanese). Each cell shows the mean score across 10 independent runs, with the standard deviation in parentheses. Rows represent the evaluator agent’s value, and columns represent the evaluated agent’s value.
@@
Table 3: Relationship between value similarity and mutual evaluation results (Trust). Means are reported by similarity level.
English Japanese
Housing Hobbies Housing Hobbies
Jonckheere–Terpstra test results Basic
Higher
-order
Basic
Higher
-order
Basic
Higher
-order
Basic
Higher
-order
JT 312999 10577 1199899 39666 2619927 89071 4564026 157416
$z$ 18.085 9.773 22.136 11.593 24.230 14.044 25.866 15.852
$p$ <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001
Table 4: Jonckheere–Terpstra test results (Trust).
@@
Table 5: Relationship between value similarity and mutual evaluation results (IOS). Means are reported by similarity level.
English Japanese
Housing Hobbies Housing Hobbies
Jonckheere–Terpstra test results Basic
Higher
-order
Basic
Higher
-order
Basic
Higher
-order
Basic
Higher
-order
JT 314945 9321 1226302 39564 2566861 82703 4422563 145235
$z$ 19.324 8.000 24.475 12.657 23.919 13.076 24.995 14.636
$p$ <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001 <.001<.001
Table 6: Jonckheere–Terpstra test results (IOS).
@@
Table 7: Pearson’s correlation coefficient between basic values and higher-order values. ($^{*}p<.05$, $^{**}p<.01$)
@@ Discussion
The results of the preliminary experiment on value controllability revealed that the perspective controllability scores varied depending on the target language model, prompt format, and language (Tables 1 and 2). First, a comparison of the performance of the models revealed that Gemini achieved the highest average scores in both English and Japanese. This suggests that Gemini has a better ability to generate responses aligned with a specified perspective than the other models under the conditions of this study. Next, we observed two trends regarding the influence of the prompt format. First, controlling the models based on the ten basic values was consistently more challenging than controlling the models based on the four higher-order values. It is conceivable that, because the latter represents broader and more abstract concepts, the models can more easily grasp them as general directives, leading to improved controllability. Second, the effects of grammatical persons varied among the models. For Gemini, second-person prompts yielded greater controllability than third-person prompts. By contrast, GPT exhibited the opposite trend, in which third-person prompts tended to result in higher controllability. Furthermore, comparing across languages, we also found that English exhibits higher controllability than Japanese, as indicated by the results of the t-test. This is likely attributable to the fact that the training data for several large language models are predominantly in English. Consequently, we infer that the models’ comprehension and adherence to instructions were higher for English than for Japanese. The existence of this linguistic bias poses a significant challenge for applying LLMs to value-based systems in non-Anglophone contexts.
@@
A further comparison between the English and Japanese results revealed a notable difference: whereas scores were stably calculated in the English experiments, the Japanese experiments exhibited a significant number of missing values, which corresponded to response refusals by the models (Tables 1 and 2). Communication in Japan is often discussed as a classic example of a high-context culture [17], and recent studies have indicated that LLMs reflect the social biases present in training data [18]. Based on these findings, it is conceivable that LLMs learn not only vocabulary and grammatical structures but also statistically internalize communication styles rooted in a high-context culture, such as avoiding direct assertions and circumventing conflict, during their training on Japanese text data. Therefore, we infer that achieving effective value control in the Japanese context requires sophisticated prompt engineering that goes beyond merely presenting values. Such prompts must also signal to the model the meta-level situational judgment such that “an explicit and assertive evaluation is necessary in this context.” Thus, our findings suggest that the value control of LLMs is a localization challenge that necessitates a deep understanding of the linguistic and cultural context in which the model operates.
@@
Subsequently, the results of the main experiment are discussed. As summarized in Tables 3, 4, 5, and 6, a higher similarity in values, based on Schwartz’s theory of basic values, was associated with higher mutual trust ratings and a greater sense of interpersonal closeness. This finding is consistent with a broad range of previous studies demonstrating the importance of value similarity in trust and relationship-building [19, 12, 11]. Therefore, our results can be interpreted as successful replication of these findings within the computational framework of mutual evaluation. Furthermore, as summarized in Table 7, a positive correlation was observed between the evaluation patterns of the ten basic values and the four higher-order values under most conditions. This finding suggests that Schwartz’s theory [6, 7] was also reproduced within the mutual evaluation framework employed in this study.
@@
Differences in the mutual evaluations based on the dialogue topic were also observed. Two-way ANOVAs with language and topic as factors for each evaluation metric (trust / IOS) and each value classification (basic / higher-order) revealed that, across all patterns, the main effect of topic was significant. Specifically, in collaborative decision-making tasks about housing, evaluations of trust and IOS were significantly higher compared to dialogues about hobbies. This finding can be explained by the relationship development process. Interpersonal relationships deepen from the initial stage of simple self-disclosure to a more advanced stage of sharing fundamental values, and the depth of the relationship is linked to trust [21, 22]. In the present study, dialogue about hobbies was assumed to mimic the superficial level of interaction characteristic of the initial stages of a relationship, whereas the cooperative decision-making task was considered to represent interaction at a deeper level, involving the sharing of values. Therefore, the difference in evaluation scores observed in this experiment suggests that topic of dialogue can simulate the process of relationship development.
@@
Furthermore, differences in the mutual evaluations based on language were observed. Except for the evaluation of trust under higher-order values control, the ANOVA results demonstrated a significant main effect of language across all other patterns. Specifically, evaluations of trust and IOS were significantly higher in the Japanese context compared to the English context. This trend suggests that the simulation reflects the tendency of the Japanese culture to emphasize harmony and relationships with others [13]. Moreover, the correlation analysis summarized in Table 7 reveals lower correlation coefficients for Japanese than for English. Specifically, no significant correlation was found between the evaluation patterns of basic and higher-order values in cooperative decision-making regarding housing tasks. This may indicate that the structure of values depends on cultural context [6, 13]. Thus, the deviation in the pattern observed under Japanese conditions in this study may reflect this cultural dependency.
Interestingly, in the higher-order values control, the effect of language disappeared, and only the main effect of topic was observed for trust evaluation. This pattern may be explained by the difference in the level of abstraction between basic and higher-order values in Schwartz’s theory. The ten basic values are relatively fine-grained and may carry culture-specific nuances, making trust judgments more sensitive to cultural signals conveyed through language. By contrast, the four higher-order values represent broader and more abstract motivational dimensions that are more universally shared across cultures [6, 23]. When agents are guided by higher-order values, the influence of language as a cultural signal may attenuate, and trust judgments may instead rely more heavily on the concrete topic of dialogue. In contrast, IOS scores showed consistent language effects across both basic and higher-order values control. This divergence indicates that trust and IOS capture partially different aspects of interpersonal evaluation. While trust is the willingness to be vulnerable to another party based on positive expectations [24], the IOS scale is a measure of closeness represented by the perceived overlap between the self and another [16]. The latter appears to remain sensitive to linguistic context even when values are framed at a more abstract level.
@@
Finally, we address the limitations of this study and outline directions for future research. While the present study employed specific dialogue tasks to simulate relationship development, future work could involve experiments under more diverse or controlled conditions to disentangle the effects of the task setting itself. Another compelling direction is to focus on the behavioral consistency and traits of individual agents. Although our work has centered on the influence of value similarity, investigating the correlation between an agent’s tendency to be a lenient or strict evaluator and the ratings it receives from others presents a valuable avenue for future research. Furthermore, a key future task will involve constructing a model that considers the interactions among multiple values. In the present experiment, the value provided to the LLM in the prompt was singular; therefore, the influence of other coexisting values could not be considered. However, humans often hold multiple values and make decisions based on trade-offs between them [25]. Therefore, modeling this value plurality is expected to enable a more realistic value simulation.
We propose the development of a more comprehensive agent model that integrates both stable base values and more flexible “prioritized” values that may shift depending on emotions and contextual reactions. Human decision-making is influenced not only by values but also by various cognitive and affective factors, such as momentary emotions and the inference of the intentions of others [26]. Incorporating these elements into the agent architecture is likely to enable more human-like simulations under realistic conditions.
In interpreting our findings, it is important to consider the choice of value framework. While the refined theory distinguishes 19 more specific values, these can be recombined into the original ten basic values [23]. Many studies still employ measures based on the ten values (e.g., [8, 9]), and we therefore adopted this framework to ensure comparability with previous research. Nevertheless, we acknowledge that the more fine-grained value model represents a promising direction for future extensions.
This study was confined to two languages, English and Japanese. This choice was theoretically motivated, as it allowed for a focused comparison between a dominant training language of LLMs (English) and a high-context language with relatively less representation in training data (Japanese). A significant direction for future research will be to expand the experimental framework to a broader range of languages and cultural contexts, thereby testing the universality of the value similarity principle in agent simulations across diverse cultural settings.
Another important limitation concerns the motivational asymmetry between humans and LLM agents. Humans are strongly motivated to evaluate trust cautiously, since misplaced trust can threaten survival, well-being, and social relationships [27, 28]. In contrast, LLMs lack such intrinsic motivations. Recent research, however, suggests promising directions for bridging this gap. Studies on agent architectures that integrate internal states and reflection [3, 29], research on intrinsic motivation in artificial intelligence [30, 31], and the use of explicit prompting strategies point toward ways of endowing agents with human-like motivational structures. Such models would enable simulations that more closely align with the human tendency to be cautious when trust carries real stakes, thereby enhancing the realism and explanatory power of LLM-based trust research.
In addition, we asked the LLM agents to make evaluations using the IOS diagrams. Previous studies have reported that multimodal LLMs can analyze images and extract their psychological meaning (e.g., [32, 33]). Moreover, the gemini-1.5-flash model used in our main experiment has been described in Google’s technical documentation as having strong visual processing capabilities [34]. At the same time, unlike human participants who may attribute rich social meaning to the IOS diagrams, the LLM’s judgments are likely constrained by its visual–textual alignment. We therefore note this as a limitation of the present study.